OpenStudy (anonymous):

A triangle has a leg that measures 6 and a hypotenuse of twelve. Which leg measure would form a right triangle?

jimthompson5910 (jim_thompson5910):

hint: a^2 + b^2 = c^2 6^2 + x^2 = 12^2 36 + x^2 = 144 Keep going to solve for x (I'll let you take over from here)